# SNF-ROM: Projection-based nonlinear reduced order modeling with smooth neural fields

### v1

@@ Abstract
Reduced order modeling lowers the computational cost of solving PDEs by learning a low-order spatial representation from data and dynamically evolving these representations using manifold projections of the governing equations. While commonly used, linear subspace reduced-order models (ROMs) are often suboptimal for problems with a slow decay of Kolmogorov $n$-width, such as advection-dominated fluid flows at high Reynolds numbers. There has been a growing interest in nonlinear ROMs that use state-of-the-art representation learning techniques to accurately capture such phenomena with fewer degrees of freedom. We propose smooth neural field ROM (SNF-ROM), a nonlinear reduced modeling framework that combines grid-free reduced representations with Galerkin projection. The SNF-ROM architecture constrains the learned ROM trajectories to a smoothly varying path, which proves beneficial in the dynamics evaluation when the reduced manifold is traversed in accordance with the governing PDEs. Furthermore, we devise robust regularization schemes to ensure the learned neural fields are smooth and differentiable. This allows us to compute physics-based dynamics of the reduced system nonintrusively with automatic differentiation and evolve the reduced system with classical time-integrators. SNF-ROM leads to fast offline training as well as enhanced accuracy and stability during the online dynamics evaluation. We demonstrate the efficacy of SNF-ROM on a range of advection-dominated linear and nonlinear PDE problems where we consistently outperform state-of-the-art ROMs.
@@
journal: Computer Methods in Applied Mechanics and Engineering
@@ 1 Introduction
Computational simulations of physical phenomena have become indispensable in scientific modeling and discovery. However, attaining a reasonable accuracy for practical problems requires resolving a wide range of spatial and temporal scales of motion, leading to large compute and memory costs. For example, the cost of direct numerical simulations, which attempt to capture all energy-containing length scales in a fluid flow, scales super-linearly with the Reynolds number of the flow, thus limiting the scope of fully resolved simulations to highly simplified problems [1, 2]. Therefore, access to high-fidelity numerical partial differential equation (PDE) solutions remains prohibitively expensive for problems requiring repeated model evaluations, such as design optimization [3] and uncertainty quantification. Reduced order models (ROMs) have been studied for the past decade [4] to alleviate the cost of model evaluations for downstream applications like control and parameter estimation. ROMs are hybrid physics and data-based methods that decouple the computation into two stages: an expensive offline stage and a cheap online stage. In the offline stage, a low-dimensional spatial representation is learned from simulation data by projecting the solution field snapshots onto a low-dimensional manifold that can faithfully capture the relevant features in the dataset. The online stage then involves evaluating the model at new parametric points by time-evolving the learned spatial representation following the governing PDE system with classical time integrators. Computational savings are achieved due to the time-evolution of reduced states with much smaller dimensionality than the full order model.
@@
Nonlinear ROMs typically use convolutional autoencoder models that limit their applicability to fixed uniform grids. This limitation is addressed by the continuous ROM (CROM) method [13, 15], which employs neural field representations [16] that can be queried anywhere in the domain. CROM decreases the cost of the online stage by solving the ROM equations at a small subset of points using hyper-reduction by over-collocation [17, 13]. However, a limitation of CROM is that the encoder module remains grid-dependent. As such, model training requires the problem to have a fixed grid structure, which may be infeasible for adaptive grid simulations and for problems where parts of the data are missing. Another major impediment to developing nonlinear ROMs is that training autoencoder networks with stochastic optimization takes much larger compute time than computing a linear ROM representation with POD. State-of-the-art continuous neural field architectures, such as CROM [13], take $O(100,000)$ epochs to train. As such, training time can be prohibitive even for moderately-sized problems.
@@
The confluence of these two features produces desirable effects on the dynamics evaluation, such as greater accuracy, robustness to hyperparameter choice, and robustness to numerical perturbations. We demonstrate the effectiveness of our method using several test cases, such as the scalar transport problem, the Burgers’ problem at high Reynolds numbers, and the Kuramoto-Sivashinsky (KS) problem. Our model implementation and several examples are available on GitHub at https://github.com/CMU-CBML/NeuralROMs.jl.
@@ 3.2 CAE-ROM: Nonlinear manifold representation
The implementation of $h^{\prime}_{\text{CAE}}$ is detailed in A.
@@ 4 SNF-ROM: Offline manifold construction
In the following subsections, we discuss the formulation of the ROM manifold in Section 4.1, and neural field regularization in Section 4.2. A comprehensive set of implementation details are presented in B, which describes the SNF-ROM projection function $h_{\theta}$, and in C, which describes the training procedure for SNF-ROM, to supplement the discussion in this section.
@@ 4.2.2 SNFW-ROM: Smooth neural field ROM with weight regularization
Compared to $L_{2}$ regularization, which penalizes both weights and biases, our weight regularization scheme leads to faster convergence and comparably smooth models. We believe this is because $L_{2}$ regularization penalizes bias vectors, which do not contribute to the high-frequency noise in the derivative. Numerical experiments in Section 6 indicate that SNFW-ROM trained with a large value of $\gamma$ consistently outperforms the other benchmark models.
@@ 5.1 Manifold projection
As illustrated in LABEL:fig:online, the initial condition $\boldsymbol{u}(0)=\boldsymbol{u}_{0}(\boldsymbol{x})$ to Eq. 1 is projected onto the manifold with $h_{\text{ROM}}$ to obtain the corresponding ROM state vector $\tilde{u}(0)$. This step is only done once to initialize the dynamics evaluation. For POD-ROMs, the projection function is a matrix-vector product as described in Eq. 19. For CAE-ROM, a preliminary approximation to $\tilde{u}(0)$ is given by the encoder prediction $e_{\theta_{e}}(\overline{u}(0))$, where $\overline{u}$ is the FOM state vector corresponding to the initial condition. Our experiments indicate that $e_{\theta_{e}}(\overline{u}(0))$ may not be the optimal projection. As such, the accuracy of the dynamics evaluation can be significantly improved by employing Gauss-Newton iteration to further optimize the return value. CAE-ROM thus solves the nonlinear system in Eq. 65 with $N_{\text{FOM}}$ equations for $N_{\text{ROM}}$ unknowns with Gauss-Newton iteration where the initial guess is taken to be the encoder prediction $e_{\theta_{e}}(\overline{u}(0))$. Further details are presented in A.
For SNF-ROM, a preliminary approximation to $\tilde{u}(0)$ is given by the intrinsic ROM manifold $\Xi_{\varrho}(0)$. Our experiments indicate that using Gauss-Newton iteration to solve Eq. 69 after the process leads to marginal improvements in the accuracy of the dynamics evaluation. As such, we select a small set of points $X_{\text{proj}}\subset\Omega$ where Eq. 66 is satisfied. The resulting system is much smaller (see Eq. 69), with |$\absolutevalue{X_{\text{proj}}}$ | equations and $N_{\text{ROM}}$ unknowns, and solved with Gauss-Newton iteration where the initial guess is given by $\Xi_{\varrho}$.
@@ 5.2 Manifold dynamics
We aim to satisfy Eq. 41 at points in $X_{\text{proj}}\subset\Omega$. For CAE-ROM, $X_{\text{proj}}=X_{\text{FOM}}$ as CAE architectures predict the entire solution field at inference time [11]. As such, the compute cost for CAE-ROM scales $\sim\mathcal{O}(N_{\text{FOM}})$ at every time step. In comparison, $X_{\text{proj}}$ can be a small subset of $\Omega$ for SNF-ROM, which leads to substantial computational savings. We restrict the semi-discretized system (Eq. 41) to $X_{\text{proj}}$ and write it in matrix-vector notation as
@@ 6 Numerical experiments
In this section, we validate the SNFL-ROM and SNFW-ROM on several advection-dominated PDE problems and compare them to POD-ROMs and CAE-ROMs [11]. Another new approach that relies on neural fields, CROM [13], was also tested. However, we found the dynamics evaluation in CROM to be sensitive to the approach used for numerical derivative computation while also diverging for reduced state vector trajectories within the training dataset. The concurrent work in [30] employs a continuous low-rank architecture to develop an alternate grid-free neural representation for reduced order modeling. Future work may consider a detailed comparison of SNF-ROM with this approach.
We use a Fourier spectral PDE solver FourierSpaces.jl [32] to generate the FOM datasets. A strong stability-preserving Runge-Kutta time integrator, available in package OrdinaryDiffEq.jl [33], is used for temporal discretization. For each experiment, $X_{\text{FOM}}$ is a uniform rectilinear grid in the problem domain, and $T_{\text{FOM}}$ are uniformly sampled from $t=0$ through $t=T$. Details of the dataset for each test case are presented in Table 1. Note that problems with intrinsic manifold dimension of $1$ are reproductive examples where the parameter $\boldsymbol{\mu}$ is held constant, implying that the ROM dynamics are evaluated on the same trajectory as the model is trained. Such examples help disambiguate errors from parameter interpolation and errors from deviations in the dynamics evaluation.
@@
$\Delta t=\frac{T}{N_{t}-1}.$ (49)
Our numerical experiments revealed no substantial improvement in performance with higher order time-integration schemes such as RK4. In all the experiments below, we set $X_{\text{proj}}$ to be the same as $X_{\text{FOM}}$ for fair comparison with the CAE-ROM. We efficiently compute Jacobian matrices by propagating dual numbers through $g_{\text{ROM}}$ in the forward mode AD procedure. Spatial derivatives for POD-ROM and CAE-ROM are computed with a second order central finite difference stencil on the $X_{\text{FOM}}$ grid. For SNFL-ROM and SNFW-ROM, spatial derivatives are computed grid-free with forward mode AD.
@@
The error in solutions predicted by ROMs is quantified as
@@
where $\boldsymbol{u}_{\text{FOM}}$ and $\boldsymbol{u}_{\text{ROM}}$ are spatiotemporal snapshots of the FOM and ROM solutions, respectively. The evolution of the error in solution predicted by ROMs is quantified using
@@
Figure 4: 1D scalar advection test case: error evolution of ROM predictions in time for dynamics evaluation with the (a) baseline $\Delta t$, and (b) 10 $\times$ $\Delta t$. The error for CAE-ROM grows with time for both cases, while the error for SNFW-ROM remains consistently low. The curve for POD-ROM is omitted from (b) due to large error values.
@@
A comparison of ROM predictions with the FOM solution is shown in Figure 3 for $t=1$ and $t=4$. The figure indicates that the nonlinear ROMs give close predictions to FOM in both instances. On the other hand, POD-ROM leads to oscillations in the solution that become more prominent near the wave. A more detailed analysis of the differences between the ROM approaches requires assessing the temporal variation of the error, as defined in Eq. 51. Figure 4 presents the error evolution of ROMs from a dynamics evaluation with the baseline $\Delta t$, and a dynamics evaluation with a 10×10$\times$10 × larger $\Delta t$. POD-ROM exhibits a large error at all times, indicating that a larger ROM representation is required to capture physics accurately. The curve for POD-ROM is omitted from Figure 4(b) due to large errors. In contrast, the nonlinear ROMs yield lower errors with a much smaller $N_{\text{ROM}}$. Despite a low initial value, the error in CAE-ROM grows steadily with time for both $\Delta t$, and reaches close to the POD-ROM error levels. The two SNF-ROM variants produce much lower errors for both $\Delta t$, though we observe a dramatic rise in error for SNFL-ROM with the larger time step size between $t=0.5$ and $t=3.5$. In contrast, SNFW-ROM produces consistently low errors for both $\Delta t$.
@@
We analyze the trajectories of the ROM state vector for the nonlinear ROMs in Figure 5 to assess deviation between predictions of $\tilde{u}(t)$ learned during training (solid lines) and $\tilde{u}(t)$ obtained from dynamics evaluations (dotted lines correspond to a dynamics evaluation with the baseline $\Delta t$, and dashed lines correspond to a dynamics evaluation with a 10×10$\times$10 × larger $\Delta t$). The first and second components of $\tilde{u}(t)=\begin{bmatrix}\tilde{u}_{1}(t)&\tilde{u}_{2}(t)\end{bmatrix}^{T}$ are represented by blue and red curves respectively. Figure 5 indicates that the three nonlinear ROMs learn ROM state trajectories that are smooth and continuous. SNFL-ROM and SNFW-ROM explicitly force this behavior by modeling $\tilde{u}$ as a simple and smooth function of $t$, which follows the expected trajectory accurately. For CAE-ROM, despite the initial proximity to the learned trajectory, ROM states from the dynamics evaluation with baseline $\Delta t$ deviate from the learned trajectory. This deviation corresponds to the growth in error with time for CAE-ROM presented in Figure 4 although further studies need to be conducted to ascertain causality. Furthermore, the ROM state trajectories for CAE-ROM obtained using a 10×10$\times$10 × larger $\Delta t$ quickly deviate from the learned trajectory. This experiment illustrates that the dynamics evaluations of SNFL-ROM and SNFW-ROM are more accurate than those of other ROMs and support taking larger time steps.
@@
Figure 12: 2D viscous Burgers test case: comparison of ROM predictions with FOM solution along the slice $y=x$. CAE-ROM, SNFL-ROM and SNFW-ROM capture the steep shock at $t=0.5$, whereas POD-ROM suffers from large oscillations near the shock.
Figure 13: 2D viscous Burgers test case: temporal variation of error in ROM solutions. SNFL-ROM and SNFW-ROM produce low errors that do not grow with time, whereas errors for POD-ROM and CAE-ROM are high and grow with time.

### v2

@@ Abstract
Reduced order modeling lowers the computational cost of solving PDEs by learning a low-dimensional spatial representation from data and dynamically evolving these representations using manifold projections of the governing equations. The commonly used linear subspace reduced-order models (ROMs) are often suboptimal for problems with a slow decay of Kolmogorov $n$-width, such as advection-dominated fluid flows at high Reynolds numbers. There has been a growing interest in nonlinear ROMs that use state-of-the-art representation learning techniques to accurately capture such phenomena with fewer degrees of freedom. We propose smooth neural field ROM (SNF-ROM), a nonlinear reduced order modeling framework that combines grid-free reduced representations with Galerkin projection. The SNF-ROM architecture constrains the learned ROM trajectories to a smoothly varying path, which proves beneficial in the dynamics evaluation when the reduced manifold is traversed in accordance with the governing PDEs. Furthermore, we devise robust regularization schemes to ensure the learned neural fields are smooth and differentiable. This allows us to compute physics-based dynamics of the reduced system nonintrusively with automatic differentiation and evolve the reduced system with classical time-integrators. SNF-ROM leads to fast offline training as well as enhanced accuracy and stability during the online dynamics evaluation. Numerical experiments reveal that SNF-ROM is able to accelerate the full-order computation by up to 199×199$\times$199 ×. We demonstrate the efficacy of SNF-ROM on a range of advection-dominated linear and nonlinear PDE problems where we consistently outperform state-of-the-art ROMs.
@@
journal: Journal of Computational Physics
@@ 1 Introduction
Computational simulations of physical phenomena have become indispensable in scientific modeling and discovery. However, attaining a reasonable accuracy for practical problems requires resolving a wide range of spatial and temporal scales of motion, leading to large compute and memory costs. For example, the cost of direct numerical simulations, which attempt to capture all energy-containing length scales in a fluid flow, scales super-linearly with the Reynolds number of the flow, thus limiting the scope of fully resolved simulations to highly simplified problems [1, 2]. Therefore, access to high-fidelity numerical partial differential equation (PDE) solutions remains prohibitively expensive for problems requiring repeated model evaluations, such as design optimization [3] and uncertainty quantification. Reduced order models (ROMs) have been studied for the past decade [4] to alleviate the cost of model evaluations for downstream applications like controls and parameter estimation. ROMs are hybrid physics and data-based methods that decouple the computation into two stages: an expensive offline stage and a cheap online stage. In the offline stage, a low-dimensional spatial representation is learned from simulation data by projecting the solution field snapshots onto a low-dimensional manifold that can faithfully capture the relevant features in the dataset. The online stage then involves evaluating the model at new parametric points by time-evolving the learned spatial representation following the governing PDE system with classical time integrators. Computational savings are achieved due to the time-evolution of reduced states with much smaller dimensionality than the full order model.
@@
Nonlinear ROMs typically use convolutional autoencoder models that limit their applicability to fixed uniform grids. Furthermore, the evaluation cost for CAE-ROMs scales with the grid size of the full-order computation [11, 12]. This significantly increases the runtime cost of a ROM, rendering the method unusable for practical applications. This limitation has been addressed in [12] which proposes a shallow masked autoencoder architecture whose inference cost does not scale with the full-order computation. Another method that addresses this limitation is the continuous ROM (CROM) [13, 15] which employs neural field representations [16] that can be queried anywhere in the domain. Both methods decrease the cost of the online stage by solving the ROM equations at a small set of collocation points. This approach is called hyper-reduction [12, 17, 13]. However, a limitation of both methods is that the encoder module remains grid-dependent and its inference cost scales with the full-order computation. As such, model training requires the problem to have a fixed grid structure, which may be infeasible for adaptive grid simulations and for problems where parts of the data are missing. Another major impediment to developing nonlinear ROMs is that training autoencoder networks with stochastic optimization takes much larger compute time than computing a linear ROM representation with POD. State-of-the-art continuous neural field architectures, such as CROM [13], take $O(100,000)$ epochs to train. As such, training time can be prohibitive even for moderately-sized problems.
@@
The confluence of these two features produces desirable effects on the dynamics evaluation, such as greater accuracy, robustness to hyperparameter choice, and robustness to numerical perturbations. With a combination of hyper-reduction and use of larger time-steps, SNF-ROM is faster than full-order computation by up to 199×199$\times$199 ×. We demonstrate the effectiveness of SNF-ROM using several test cases, such as the scalar transport problem, the Burgers’ problem at high Reynolds numbers, and the Kuramoto-Sivashinsky (KS) problem. Our model implementation and several examples are available on GitHub at https://github.com/CMU-CBML/NeuralROMs.jl.
@@ 3.2 CAE-ROM: Nonlinear manifold representation
The implementation of $h^{\prime}_{\text{CAE}}$ is detailed in A.1.
@@ 4 SNF-ROM: Offline manifold construction
In the following subsections, we discuss the formulation of the ROM manifold in Section 4.1, and neural field regularization in Section 4.2. A comprehensive set of implementation details are presented in A.2, which describes the SNF-ROM projection function $h_{\theta}$, and in C, which describes the training procedure for SNF-ROM, to supplement the discussion in this section.
@@ 4.2.2 SNFW-ROM: Smooth neural field ROM with weight regularization
Compared to $L_{2}$ regularization, which penalizes both weights and biases, our weight regularization scheme leads to faster convergence and comparably smooth models. We believe this is because $L_{2}$ regularization penalizes bias vectors, which do not contribute to the high-frequency noise in the derivative. Numerical experiments in Section 6 indicate that SNFW-ROM trained with a large value of $\gamma$ consistently outperforms the other benchmark models. A comparison of regularization approaches is presented in B.
@@ 5.1 Manifold projection
As illustrated in LABEL:fig:online, the initial condition $\boldsymbol{u}(0)=\boldsymbol{u}_{0}(\boldsymbol{x})$ to Eq. 1 is projected onto the manifold with $h_{\text{ROM}}$ to obtain the corresponding ROM state vector $\tilde{u}(0)$. This step is only done once to initialize the dynamics evaluation. For POD-ROMs, the projection function is a matrix-vector product as described in Eq. 19. For CAE-ROM, a preliminary approximation to $\tilde{u}(0)$ is given by the encoder prediction $e_{\theta_{e}}(\overline{u}(0))$, where $\overline{u}$ is the FOM state vector corresponding to the initial condition. Our experiments indicate that $e_{\theta_{e}}(\overline{u}(0))$ may not be the optimal projection. As such, the accuracy of the dynamics evaluation can be significantly improved by employing Gauss-Newton iteration to further optimize the return value. CAE-ROM thus solves the nonlinear system in Eq. 66 with $N_{\text{FOM}}$ equations for $N_{\text{ROM}}$ unknowns with Gauss-Newton iteration where the initial guess is taken to be the encoder prediction $e_{\theta_{e}}(\overline{u}(0))$. Further details are presented in A.1.
For SNF-ROM, a preliminary approximation to $\tilde{u}(0)$ is given by the intrinsic ROM manifold $\Xi_{\varrho}(0)$. Our experiments indicate that using Gauss-Newton iteration to solve Eq. 70 after the process leads to marginal improvements in the accuracy of the dynamics evaluation. As such, we select a small set of points $X_{\text{proj}}\subset\Omega$ where Eq. 67 is satisfied. The choice of points in $X_{\text{proj}}$ is not restricted to $X_{\text{FOM}}$ and can be sampled anywhere in $\Omega$. Although only $N_{\text{ROM}}$ points are needed for solving Eq. 68, the system is typically over-determined [13] with $N_{\text{ROM}}<\absolutevalue{X_{\text{proj}}}<<N_{\text{FOM}}$ where |$\absolutevalue{X_{\text{proj}}}$ | is the number of points in $X_{\text{proj}}$. The resulting system is much smaller (see Eq. 70), with |$\absolutevalue{X_{\text{proj}}}$ | equations and $N_{\text{ROM}}$ unknowns, and solved with Gauss-Newton iteration where the initial guess is given by $\Xi_{\varrho}$.
@@ 5.2 Manifold dynamics
We aim to satisfy Eq. 41 at points in $X_{\text{proj}}\subset\Omega$. As such, we refer to $X_{\text{proj}}$ as the set of hyper-reduction collocation points. For CAE-ROM, $X_{\text{proj}}=X_{\text{FOM}}$ as CAE architectures do not implement hyper-reduction, and predict the entire solution field at inference time [11]. As such, the cost for CAE-ROM scales with that of the FOM. In comparison, $X_{\text{proj}}$ can be a small subset of $\Omega$ for SNF-ROM, which leads to substantial computational savings. We restrict the semi-discretized system (Eq. 41) to $X_{\text{proj}}$ and write it in matrix-vector notation as
@@
An alternative time-evolution strategy, called least-squares Petrov-Galerkin (LSPG) [31], is obtained by applying a time-discretization to Eq. 43. The problem is then evolved by solving the nonlinear system for $\tilde{u}$ at the next time-step with nonlinear least squares. We consider explicit time-integrators in this paper, for which LSPG has been shown to be equivalent to Galerkin projection [31]. As such, we focus only on Galerkin projection method for the rest of this paper.
@@ 6 Numerical experiments
In this section, we validate the SNFL-ROM and SNFW-ROM on several advection-dominated PDE problems and compare them to POD-ROMs and CAE-ROMs [11]. The concurrent work in [30] employs a continuous low-rank architecture to develop an alternate grid-free neural representation for reduced order modeling. Future work may consider a detailed comparison of SNF-ROM with this approach.
We use a Fourier spectral PDE solver FourierSpaces.jl [32] to generate the FOM datasets. A strong stability-preserving Runge-Kutta time integrator, available in package OrdinaryDiffEq.jl [33], is used for temporal discretization. For each experiment, $X_{\text{FOM}}$ is a uniform rectilinear grid in the problem domain, and $T_{\text{FOM}}$ are uniformly sampled from $t=0$ through $t=T$. Details of the dataset for each test case are presented in Table 1. Note that problems with an intrinsic manifold dimension of $1$ are reproductive examples where the parameter $\boldsymbol{\mu}$ is kept constant, implying that the ROM dynamics are evaluated on the same trajectory as the model is trained. Such examples help to disambiguate errors from parameter interpolation and errors from deviations in the dynamics evaluation.
@@
$\Delta t_{0}=\frac{T}{N_{t}-1}.$ (49)
Our numerical experiments revealed no substantial improvement in performance with higher-order time-integration schemes such as RK4. The accuracy of all ROMs deteriorates with the time-step size. Our experiments in this section reveal that the accuracy of SNF-ROM does not significantly degrade with increasing $\Delta t$. Taking large time-steps can thus lead to a substantial speedup without compromising accuracy of the method.
[…]
We efficiently compute Jacobian matrices by propagating dual numbers through $g_{\text{ROM}}$ in the forward mode AD procedure. As such, the cost of computing $\mathbf{J}_{g}(\tilde{u};X_{\text{proj}})$ is only $N_{\text{ROM}}$ times that of a single forward pass, $g_{\text{ROM}}$, evaluated on $X_{\text{proj}}$. This cost is independent of the FOM computation when hyper-reduction is employed. Spatial derivatives for POD-ROM and CAE-ROM are computed with a second order central finite difference stencil on the $X_{\text{FOM}}$ grid. For SNFL-ROM and SNFW-ROM, spatial derivatives are computed grid-free with forward mode AD.
@@
We quantify the projection error for each ROM method as
$e_{\text{proj}}$ $=\left(\frac{1}{\absolutevalue{T_{\text{FOM}}}\absolutevalue{M_{\text{ROM}}}}\sum_{\begin{subarray}{c}t\in T_{\text{FOM}}\\ \boldsymbol{\mu}\in M_{\text{FOM}}\end{subarray}}\frac{\norm{\overline{u}(t;\boldsymbol{\mu})-g_{\text{ROM}}\circ h_{\text{ROM}}(\overline{u}(t;\boldsymbol{\mu}))}_{2}^{2}}{\norm{\overline{u}(t;\boldsymbol{\mu})}_{2}^{2}}\right)^{1/2}.$ (50)
This metric assesses the ability of a model order reduction approach to capture the solution manifold for a PDE problem. The relative error as a function of $(\boldsymbol{x},t;\boldsymbol{\mu})$ in ROM predictions is quantified as
@@
where $\boldsymbol{u}_{\text{FOM}}$ and $\boldsymbol{u}_{\text{ROM}}$ are spatiotemporal snapshots of the FOM and ROM solutions, respectively. The evolution of relative error in ROM predictions is quantified using
@@
Figure 4: 1D scalar advection test case: comparison of projection error $e_{\text{proj}}$ of different ROMs as a function of ROM dimension $N_{\text{ROM}}$. The three nonlinear ROMs adequately capture the solution manifold with $N_{\text{ROM}}=1$ which is equal to the intrinsic manifold dimension of the problem. Comparatively, the linear POD-ROM requires $N_{\text{ROM}}=16$ modes to capture the solution manifold for this case.
Figure 5: 1D scalar advection test case: error evolution of ROM predictions in time for dynamics evaluation with (a) $\Delta t=\Delta t_{0}$, and (b) $\Delta t=10\Delta t_{0}$. The error for CAE-ROM grows with time for both cases, while the error for SNFW-ROM remains consistently low. The curve for POD-ROM is omitted from (b) due to large error values.
@@
A comparison of ROM predictions with the FOM solution is shown in Figure 3 for $t=1$ and $t=4$. The figure indicates that the nonlinear ROMs give close predictions to FOM in both instances. On the other hand, POD-ROM leads to oscillations in the solution that become more prominent near the wave.
To assess the model order reduction capability of different ROMs, we compute the projection error as a function of ROM dimension $N_{\text{ROM}}$ in Figure 4. As expected, the linear POD-ROM requires a large number of modes to capture the solution manifold. This is because advection dominated problems exhibit a slow decay in their Kolmogorov $n$-width. In contrast, the three nonlinear ROMs are able to adequately capture the solution manifold with $N_{\text{ROM}}=1$, which is the intrinsic manifold dimension of this problem. Specifically, SNFW-ROM best captures the solution manifold for small $N_{\text{ROM}}$. We also note that increasing $N_{\text{ROM}}$ beyond the intrinsic manifold dimension of the problem does not lead to a significant improvement in the performance of nonlinear models.
@@
A more detailed analysis of the differences between the ROM approaches requires assessing the temporal variation of the error, as defined in Eq. 52. Figure 5 presents the error evolution of ROMs from a dynamics evaluation with $\Delta t=\Delta t_{0}$, and a dynamics evaluation with a $\Delta t=10\Delta t_{0}$. POD-ROM exhibits a large error at all times, indicating that a larger ROM representation is required to capture physics accurately. The curve for POD-ROM is omitted from Figure 5(b) due to large errors. In contrast, the nonlinear ROMs yield lower errors with a much smaller $N_{\text{ROM}}$. Despite a low initial value, the error in CAE-ROM grows steadily with time for both $\Delta t$, and reaches close to the POD-ROM error levels. The two SNF-ROM variants produce much lower errors for both $\Delta t$, though we observe a dramatic rise in error for SNFL-ROM with the larger time step size between $t=0.5$ and $t=3.5$. In contrast, SNFW-ROM produces consistently low errors for both $\Delta t$.
@@
We analyze the trajectories of the ROM state vector for the nonlinear ROMs in Figure 6 to assess deviation between predictions of $\tilde{u}(t)$ learned during training (solid lines) and $\tilde{u}(t)$ obtained from dynamics evaluations (dotted lines correspond to a dynamics evaluation with $\Delta t=\Delta t_{0}$, and dashed lines correspond to a dynamics evaluation with $\Delta t=10\Delta t$). The first and second components of $\tilde{u}(t)=\begin{bmatrix}\tilde{u}_{1}(t)&\tilde{u}_{2}(t)\end{bmatrix}^{T}$ are represented by blue and red curves respectively. Figure 6 indicates that the three nonlinear ROMs learn ROM state trajectories that are smooth and continuous. SNFL-ROM and SNFW-ROM explicitly force this behavior by modeling $\tilde{u}$ as a simple and smooth function of $t$, which follows the expected trajectory accurately. For CAE-ROM, despite the initial proximity to the learned trajectory, ROM states from the dynamics evaluation with $\Delta t=\Delta t_{0}$ deviate from the learned trajectory. This deviation corresponds to the growth in error with time for CAE-ROM presented in Figure 5 although further studies need to be conducted to ascertain causality. Furthermore, the ROM state trajectories for CAE-ROM obtained using a 10×10$\times$10 × larger $\Delta t$ quickly deviate from the learned trajectory. This experiment illustrates that the dynamics evaluations of SNFL-ROM and SNFW-ROM are more accurate than those of other ROMs and support taking larger time steps.
6.2 2D scalar advection test case
@@
[…]
                         Number of hyper-reduction points (|$(\absolutevalue{X_{\text{proj}}})$
$16,384$ $4,096$ $1,024$ $256$ $64$
[…]
The memory allocations diminish proportionally as $X_{\text{proj}}$ is reduced and $\Delta t$ is increased. This leads to a consistent reduction in wall-time till |$\absolutevalue{X_{\text{proj}}}=256$. For smaller array sizes in this range, the wall time is dominated by the latency in communication between the host CPU and the GPU device. For that reason, we do not see a substantial increase in speed-up between |$\absolutevalue{X_{\text{proj}}}=256$ and |$\absolutevalue{X_{\text{proj}}}=64$.
We consider the variation of relative error with $\Delta t$ and |$\absolutevalue{X_{\text{proj}}}$ |. In this experiment, we note that relative error does not significantly vary with |$\absolutevalue{X_{\text{proj}}}$ |. However, the error increases substantially as we increase the time step size $\Delta t$. Note that in this work, $X_{\text{proj}}$ is obtained by uniformly coarsening the FOM grid. A more sophisticated strategy for choosing $X_{\text{proj}}$ may further lower error for the same |$\absolutevalue{X_{\text{proj}}}$ |. Finally, we note the following oddity: the relative error goes down as we reduce |$\absolutevalue{X_{\text{proj}}}$ |. Further investigation is necessary to ascertain the reason behind this trend.
@@
Figure 13: 2D viscous Burgers test case evalauted at $\mu=1.00$ (interpolation): comparison of ROM predictions with FOM solution along the slice $y=x$. CAE-ROM, SNFL-ROM and SNFW-ROM capture the steep shock at $t=0.5$, whereas POD-ROM suffers from large oscillations near the shock.
Figure 14: 2D viscous Burgers test case evaluated at $\mu=1.00$ (interpolation): temporal variation of error in ROM solutions. SNFL-ROM and SNFW-ROM produce low errors that do not grow with time, whereas errors for POD-ROM and CAE-ROM are high and grow with time.